Circular No. 8101

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

GRB 030329 G. R. Ricker, Massachusetts Institute of Technology; and the HETE

Science Team report the discovery with the HETE satellite (+ SXC camera) of an extremely bright γ-ray burst (GRB) on Mar. 29.484 UT located at α = 10^{h}44^{m}50^{s}, δ = +21^{o}30′.9 (equinox 2000.0; position uncertainty < 2’). The GRB fluence in the band 30–400 keV was ∼ 1.2 × 10^{−4} erg cm^{−2} (i.e., ∼ 4000 Crab-s), integrated over the ∼ 50-s burst duration. This is the brightest GRB ever detected by HETE and ranks among the brightest 1 percent of all GRBs previously detected (cf. Paciesas et al. 1999, Ap.J. Suppl. 122, 465). Searches for both contemporaneous and new observations of the field of GRB 030329 (ecliptic latitude +12^{o}.5), in all wavelength bands, are encouraged.

SUPERNOVAE 2003cj, 2003ck, 2003cl, 2003cm, 2003cn Several discoveries of apparent supernovae have been reported from

unfiltered CCD images — SNe 2003cj, 2003ck, and 2003cl by W. M. Wood-Vasey, G. Aldering, and P. Nugent (NEAT; cf. IAUC 8095), and SNe 2003cm in UGC 10590 and 2003cn in IC 849 by M. Papenkova and W. Li (LOTOSS/KAIT; cf. IAUC 8097).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003cj Mar. 26.28 12^{h}15^{m}51.^{s}90 + 8^{o}59′15′′.7 20.2 0′′.8 W, 2′′.6 S
2003ck Mar. 26.39 13 06 21.72 + 3 04 21.0 19.3 3′′.1 W, 0′′.6 N
2003cl Mar. 27.24 12 38 50.11 − 2 49 58.8 19.5 2′′.2 E, 1′′.8 N
2003cm Mar. 27.5 16 50 55.02 +59 43 18.1 19.4 8′′.9 W, 10′′.9 N
2003cn Mar. 29.4 13 07 37.05 − 0 56 49.9 18.1 24′′.2 W, 16′′.0 S

Additional reported approximate magnitudes: SN 2003cj, 2002 Apr. 14.21 UT, [21.2; 2003 Feb. 12.51, [20.2; Mar. 16.36, [19.3; 21.64, [18.7; 22.47, [20.0; 27.20, 20.0. SN 2003ck, 2002 Apr. 13.2, [20.9; 2003 Feb. 3.61, [20.5; 17.45, [19.2; Mar. 28.26, 19.5. SN 2003cl, 2002 July 21.17, [20.6; 2003 Feb. 16.39, [18.6; Mar. 6.54, 20.0. SN 2003cm, 2002 July 27.3, [20; 2003 Mar. 28.5, 19.4; 29.5, 19.4. SN 2003cn, Mar. 7.4, [19.5; 30.4, 18.1.

SUPERNOVA 2003cg IN NGC 3169 An unfiltered CCD image taken on Mar. 19.511 UT by Y. Hirose,

Chigasaki, Kanagawa, Japan, shows nothing at the position of SN 2003cg (limiting mag ∼ 16.0 in poor conditions; communicated via H. Yamaoka).

2003 March 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green